Given a slab of material with area of 1.0 m2

and thickness 2.0 x 10-2 m, (a) what is 

the thermal resistance if the material is asbestos? (b) what is the thermal resistance 

if the material is iron? (c) what is the thermal resistance if the material is copper?



1
Expert's answer
2021-04-23T11:08:52-0400
"R=\frac{d}{kA}"

a)


"R=\frac{0.02}{(0.17)(1)}=0.118\frac{K}{W}"

b)


"R=\frac{0.02}{(80.2)(1)}=2.49\cdot10^{-4}\frac{K}{W}"

c)


"R=\frac{0.02}{(401)(1)}=4.99\cdot10^{-5}\frac{K}{W}"